(a) Calculate the momentum of a 1950 kg elephant charging a hunter at a speed of 7.50 m/s.
_kg·m/s

(b) Compare the elephant's momentum with that of a 0.0400 kg bullet fired at a speed of 600 m/s.
(momentum of elephant / momentum of bullet)

(c) What is the momentum of the 90.0 kg hunter running at 7.50 m/s after missing the elephant?
_ kg·m/s
Answer #1

a)p=m*v

where P is the momentum

m is the mass

V is the velocity

momentum of elephant=1950*7.5=14625 kg m/s

b) momentum of bullet=mass of bullet*velocity of bullet

p=0.04*600=24 kg m/s

Momentum of elephant is 609.375 times that of bullet

c) momentum of Hunter=mass of Hunter*velocity of Hunter

p=90*7.5=675 kg m/s